OpenStudy (casadie1004):

Fan and medal. X^2-9=0 A. X=+3 B. X=+9 C. X=0 D. No solution Can u show the steps on how to do it?

OpenStudy (bobo-i-bo):

Do you know how to rearrange equations?

OpenStudy (jdoe0001):

hint: \(\textit{difference of squares} \\ \quad \\ (a-b)(a+b) = a^2-b^2\qquad \qquad a^2-b^2 = (a-b)(a+b)\)

OpenStudy (jdoe0001):

\(\bf x^2-9=0\qquad {\color{brown}{ 9=3^2}}\implies x^2-{\color{brown}{ 3^2}}=0\)

OpenStudy (casadie1004):

i know all that but im wondering if the answer is 3 or 0

Directrix (directrix):

X^2-9=0 (x + 3) * (x - 3) = 0 x + 3 = 0 OR x - 3 = 0 x = - 3 or x = + 3 @Casadie1004

OpenStudy (bobo-i-bo):

Lol, if we're just giving answers away then: \[x^2-9=0\] \[x^2-9+9=0+9\] \[x^2=9\] \[\sqrt{x^2}=\sqrt{9}\] \[x=\pm3\]